OpenStudy (anonymous):

solve for x: -x^6 - x^3 + 12 = 0

hero (hero):

x^6 + x^3 - 12 = 0 x^3 = y y^2 + y - 12 = 0 y^2 + 4y - 3y - 12 = 0 y(y+4)-3(y+4) = 0 (y+4)(y-3) = 0 replace y = x^3.. x^3 + 4 = 0 x^3 - 3 = 0 x^3 = -4 x^3 = 3 x = cuberoot(-4) x = cuberoot(3)